What is the size of each interior angle of a regular polygon if the number of sides is 4? 5? 6? 7? 8? 9?
Answer by Alan3354(21583)

What is the size of each interior angle of a regular polygon if the number of sides is 4? 5? 6? 7? 8? 9?
-------------------
The number of sides is the same as the number of angles and vertices.
The total of the interior angles of a polygon is 180*(n-2), where n is the number of sides, angles and vertices.
For regular polygons, all angles are equal, so they're the total/n.
That is 180*(n-2)/n
Just sub the number for n:
For 4, it's 360/4 = 90 degs
For 5, it's 540/5 = 108 degs
-------------
For 9, it's 1260/9 = 140 degs
For 36, it's 6120/36 = 170 degs
Notice that the angle increase as the number of sides increase. You can use that as a check for polygons with 5, 6, 7 & 8 sides.
---------
For 1000, it's 179.64 degs
As n approaches infinity (which is a circle), the angle approaches 180 degs.